Nam Kading (01.01.2012 to 03.31.2012)

You are here

PDF versionPDF version
Year: 
2012
Quarter: 
January - March
Communications activities: 

 No significant communications activities to report

Unusual Events at the TEAM Site: 

 No unusual events to report

New Species: 

 No new species to report

Protocol Activities: 

 Terrestrial vertebrate monitoring

  • 8 January 2012 - 23 January 1012 - Set up 13 CT at 30CT at Phou Phapaek site.
  • 8 February 2012 - 23 February 3012 - Collect 13 CT from Phou Phapaek site.
  • 5 March 2012 - 23 March 2012 - Set up 17 CT at Naman site

Vegetation monitoring

  • 6 January 2012 - 25 January 2012 - Plot 6 (Talabut 2) 528 stems measured
  • 6 February 2012 - 26 February 2012 - Plot 5  (Talabut 1) 514 stems measured
  • 5 March 2012 - 21 March 2012 - Plot 1 (Naman 1) 529 stems measured
  • 22 March 2012 - 12 April 2012 - plot 2 (Naman 2) 570 stems measuredClimate monitoring

Climate monitoring

  • Climate data recieved and uploaded monthly
  • Regular maintenance carried out as per protocols.
  •  January 2012 - RH sensor changed and sent away for calibration.
Protocol Problems: 
  • No problems experienced 
Schedule Problems: 

 No problems experienced

Logistical Problems: 

One camera trap presumed stollen from the field at the Phou Talabat site (discovered during pick up). Cable had been cut through  with macheti. Equipment records have been updated according to protocols.

Suggestions: 

 No new suggestions